# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ 2 Lower Bound of Sample Complexity, for a Positive Instance
For (8), we notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,0}^{2}}$, further $T_{j}(a)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}$. That means for $n\leq T_{j}(a)$, we have $\mathcal{E}_{j}\Rightarrow\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Rightarrow\{\tau_{n}^{j}(a)<\tau\}$, which suggests (8).
@@
For (9), we introduce an alternative instance $\nu_{j,a}$, with $r_{a^{\prime}}^{\nu_{j,a}}=\begin{cases}r_{a}^{\nu}&a^{\prime}=j\\ r_{j}^{\nu}&a^{\prime}=a\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. $\nu_{j,a}$ swaps the position of arm $a$ and $j$ compared to the instance $\nu$. Denote $\tilde{\mathcal{E}}^{aj}(n)=\{\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{j}(n)=\{A_{\tau_{n}^{j}(a)+1}=j,\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{a}(n)=\{A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau\}$. Notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,a}^{2}}$, further $T_{j}(a)\leq\frac{O(1)}{\Delta_{j,a}^{2}(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})})}$, we can use Lemma 7 and 10, which apply the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), to conclude
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (10)
@@
where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$. To complete this step, we have to put $1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}$ in the denominator of $T_{j}(a)$. (10) suggests that the algorithm is unable to distinguish $\nu$ and $\nu_{ja}$ at the round $t$ such that $N_{a}(t)+N_{j}(t)\leq T_{j}(a)$. For a symmetric algorithm, we have $\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))$, which means
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (11)
By the Pinsker’s Inequality $kl(x,y)\geq 2(x-y)^{2}$ and $\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j})+\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a})=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj})$, we can derive
$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a).$
@@
The above inequality suggests that we we have proved (9), further the correctness of (C5). ∎
3 Upper Bound on the Complexity of 1-identification
@@
In this section, we present our proposed Algorithm PSEEB with rough analysis.
3.1 Algorithm
@@
Randomly generate a permutation of $\sigma:[K]\rightarrow[K]$, take $B_{b}=\{\sigma(i)\}_{i=1}^{\min\{2^{b-1},K\}}$ for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$;
// $\sigma(i)$ denotes the arm index in position $i$
@@
4Create algorithm copies $\{\text{alg}_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ by specifying $\text{alg}_{b}$ through calling Algorithm 2 with $B\leftarrow B_{b}$, $\mu_{0}\leftarrow\mu_{0}$, $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$, $C\leftarrow 1.01$, $K\leftarrow K$. Samples across bracket are not shared;
@@ 7      Execute $\text{alg}_{b}$ for one time round;
We propose Algorithm 1 to solve the 1-identification problem, which merges and adapts the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025). Katz-Samuels and Jamieson (2020) create each bracket by a unique and independent random shuffle, and a corresponding algorithm copy to pull arms in this bracket. In contrast to them, we specify all the brackets from a single random permutation. Algorithm 1 generates a random permutation of $[K]$ before the start of the pulling process. Based on the random permutation, Algorithm 1 create $\lceil\log_{2}K\rceil+1$ brackets $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$, satisfying $B_{1}\subset B_{2}\subset\cdots\subset B_{\lceil\log_{2}K\rceil+1}=[K]$. For a bracket index $1\leq b\leq\lceil\log_{2}K\rceil$, we have $|B_{b}|=2^{b}$, and $|B_{\lceil\log_{2}K\rceil+1}|=\min\{2^{\lceil\log_{2}K\rceil},K\}$. Our approach is to guarantee $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ follow an increasing order, which is crucial to our Lemma 31. Lemma 31 asserts that with at least some probability, the first several brackets contain at least 1 qualified arm, if there are multiple qualified arms.
@@
Then, Algorithm 1 calls Algorithm 2, defined in Appendix C, as an oracle to set up $\lceil\log_{2}K\rceil+1$ algorithm copies. Algorithm 2 is adapted from the Algorithm 2,3,4,5 in Li and Cheung (2025), named as Sequential Exploration Exploitation(SEE). Algorithm copy $b\in[\lceil\log_{2}K\rceil+1]$ takes bracket $B_{b}$ as an input parameter and only pull arms in $B_{b}$. To fit in the bracketing idea, we modify its budget allocation and stopping rules. Detailed discussion of Algorithm 2 is in Appendix C.
Algorithm 1 manages all the algorithm copies by a round-robin manner. In each round, Algorithm 1 execute one of the algorithm copies with one arm pull and check whether the copy terminates with an output $\hat{a}\in[K]\cup\{\textsf{None}\}$. By the design of Algorithm 2, $\text{alg}_{\lceil\log_{2}K\rceil+1}$ might terminate with taking output $\hat{a}\in[K]\cup\{\textsf{None}\}$. While for bracket $b$ whose size is smaller than $K$, $\text{alg}_{b}$ cannot output None but keep pulling arms, even if the $\text{alg}_{b}$ finds that $\max_{a\in B_{b}}\mu_{a}<\mu_{0}$. The output strategy differs with the algorithm SEE in Li and Cheung (2025).
Once there exists an algorithm copy terminates with an output, Algorithm 1 will adopt the output as the final output and terminate the overall sampling process. We require each algorithm copy to pull arms in a parallel manner without sharing samples. Denote $\tau$ as the total length of algorithm 1, and $\tau(B_{b})$ as the total length of copy $\text{alg}_{b}$. We allow $\tau,\tau(B_{b})$ to take infinities. Since any termination of algorithm copy results in the termination of Algorithm 1, we have $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$, further $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{b})$.
Generally speaking, for a positive instance $\nu$ with $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\mu_{K}$, we wish a bracket $B$ with size smaller than $K$ contains a qualified arm $j\in[m]$. Then, we wish the algorithm copy on this bracket $B$ is able to spend less pulling times to output arm $j$. While for a negative instance $\nu$ with $\mu_{0}>\mu_{1}\geq\mu_{2}\geq\cdots\mu_{K}$, it is possible that all the algorithm copies whose index $b\leq\lceil\log_{2}K\rceil$ never terminates. We only wish the copy $\lceil\log_{2}K\rceil+1$ is able to terminate with finite sampling and output $\hat{a}=\textsf{None}$ when encountering a negative instance.
3.2 Performance Guarantee of Algorithm 1
In the following of this section, we present the sketch proof of Theorem 2 and 3. We use Figure 1 in Appendix C.3 to summarize the overall steps.
Sketch Proof of Theorem 2.
The first step is prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)=0$, whose main idea is very similar to the proof of Theorem 5.2 in Li and Cheung (2025). Same as Li and Cheung (2025), denote $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ as the minimum phase index such that the concentration events of copy $b$ hold, we can prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)\leq\Pr(\max\{\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1},\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}\}=+\infty)=0$.
Given the first step, we are able to conclude $\Pr\Big(\tau<+\infty\Big)=1$, by the fact that $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Then, to prove Algorithm 1 is $\delta$-PAC, we suffice to show $\Pr_{\nu}\left(\hat{a}=\textsf{None}\text{ or }\{\hat{a}\in[K],\mu_{\hat{a}}<\mu_{0}\}\right)<\delta$ for a positive instance $\nu$, and $\Pr_{\nu}\left(\hat{a}\in[K]\right)<\delta$ for a negative instance $\nu$.
The proof for both targets is similar. Recall we deliver $\frac{\delta}{\lceil\log_{2}K\rceil+1}$ to each algorithm copy in the Line 1 o Algorithm 1. We are able to prove $\Pr(\text{alg}_{b}\text{ outputs incorrect }\hat{a})<\frac{O(1)\delta}{\lceil\log_{2}K\rceil+1}$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Taking the union bound, we complete the proof. ∎
Sketch Proof of Theorem 3.
We split the proof into two cases, which are positive and negative $\nu$.
If the instance $\nu$ is negative, we focus on analyzing $\tau(B_{\lceil\log_{2}K\rceil+1})$, which is very similar to the negative case of Theorem 5.3 in Li and Cheung (2025). By the fact that $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{\lceil\log_{2}K\rceil+1})$, we complete the proof.
If the instance $\nu$ is positive, denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$. Denote $\tilde{T}_{j}(a)=\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, $\tilde{T}_{j}(a;k)=\frac{\log\frac{4K}{\delta_{k}}+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, we are going to show $\mathbb{E}\tau\leq O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\tilde{T}_{j}(a)\right)$ holds for all $j\in[m]$. What’s more, the constant hidden by the Big O notation is independent of $j$. For any given $j\in[m]$, we introduce $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$, which is the minimum bracket index that contains at least one arm not worse than $\mu_{j}$. As we adopt a random permutation to create brackets, $b_{j}$ is a random variable, and we have $\mathbb{E}\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\mathbb{E}\tau\mathds{1}(b_{j}=b)$. Then, we turn to analyze $O(\log K)\mathbb{E}\tau(B_{b})\mathds{1}(b_{j}=b)$, which is the upper bound of $\mathbb{E}\tau\mathds{1}(b_{j}=b)$.
Denote $\tau^{\text{ee}}(B_{b}),\tau^{\text{et}}(B_{b})$ as the total length of exploration and exploitation periods of algorithm copy $\text{alg}_{b}$ separately. For any arm $a\in B_{b}$, denote $N_{a}^{\text{ee},B_{b}}$ as the pulling times of arm $a$ in the exploration period. Not hard to see $\tau(B_{b})=\tau^{\text{ee}}(B_{b})+\tau^{\text{et}}(B_{b})$. We are going to prove
If (12)(13) are both correct, we can sum up (12)(13) for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$. Lemma 31, 33, 34 prove upper bounds for $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}|B_{b}|\Pr(b_{j}=b)$ and $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}$, which complete the proof of Theorem 3.
To prove (12), we first show the algorithm must terminate no later than the end of phase $L_{b}^{\text{pos}}=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$, if $b_{j}=b$. Notice that $\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}=O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$, and Lemma 17 shows $\Pr(\kappa^{\text{ee}}_{b}\geq k)\leq\frac{\pi^{2}\delta_{k-1}}{6}$, $\Pr(\kappa^{\text{et}}_{b}\geq k)\leq\frac{\pi^{2}\delta}{6\alpha_{k-1}}$, we can take expectation on both sides of $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}\mathds{1}(b_{j}=b)$ to complete the proof of (12).
To prove (13), we first show for each $a\in B_{b}$, we have
The intuition is as follows. Given the fact that $j\in B_{b}$, once the pulling times of arm $a$ is higher than $O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}$, copy $\text{alg}_{b}$ can assert either $\text{UCB}_{a}<\mu_{j}$ or $\text{LCB}_{a}>\mu_{0}$, with a constant confidence level. In both cases, $\text{alg}_{b}$ stops pulling arm $a$ in the exploration period. By the facts that $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)=(|Q|+\sum_{a\in B_{b}}N_{a}^{\text{ee},B_{b}})\mathds{1}(b_{j}=b)$ and elements in $B_{b_{j}}$ are independent of $\kappa^{\text{ee}}_{b_{j}},\kappa^{\text{et}}_{b_{j}}$, we can sum up (14) for $a\in B_{b}$ and derive
Since $\Pr(a\in B_{b},b_{j}=b)\leq\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}$, and $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}$, we can conclude (13). ∎

### v2

@@ 2 Lower Bound of Sample Complexity, for Positive Instances
For (8), we notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,0}^{2}}$, and that $T_{j}(a)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}$. That means for $n\leq T_{j}(a)$, we have $\mathcal{E}_{j}\Rightarrow\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Rightarrow\{\tau_{n}^{j}(a)<\tau\}$, which suggests (8).
@@
For (9), we introduce an alternative instance $\nu_{j,a}$, with $r_{a^{\prime}}^{\nu_{j,a}}=\begin{cases}r_{a}^{\nu}&a^{\prime}=j\\ r_{j}^{\nu}&a^{\prime}=a\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. The instance distribution $\nu_{j,a}$ swaps the position of arm $a$ and $j$ compared to the instance $\nu$. Notice that
$\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,a}^{2}},\quad T_{j}(a)\leq\frac{O(1)}{\Delta_{j,a}^{2}(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})})}$ (10)
Lemma 8 and 11 combine (10) and the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), to conclude
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1$ (11)
@@
where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$. $1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}$ in the denominator of $T_{j}(a)$ is crucial to show (11). (11) suggests that the algorithm is unable to distinguish $\nu$ and $\nu_{ja}$ at the round $t$ such that $N_{a}(t)+N_{j}(t)\leq T_{j}(a)$. For a symmetric algorithm, we have $\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))$, which means
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1$ (12)
By the Pinsker’s Inequality $\text{kl}(x,y)\geq 2(x-y)^{2}$ and $\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n))+\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n))$, we can derive
$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1.$
@@
The above inequality suggests that we have proved (9), hence (C5) is justified. ∎
3 Algorithm Design for the Upper Bound of the Sample Complexity
@@
In this section, we present our proposed algorithm, Parallel Sequential Exploration Exploitation on Brackets (PSEEB), together with a high-level discussion of its design and analysis.
@@
Randomly generate a permutation of $\sigma:[K]\rightarrow[K]$, take $B_{b}=\{\sigma^{-1}(i)\}_{i=1}^{\min\{2^{b-1},K\}}$ for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$;
// $\sigma^{-1}(i)$ denotes the arm index in position $i$
@@
4Create algorithm copies $\{\text{alg}_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ by specifying $\text{alg}_{b}$ through calling Algorithm 2 with $B\leftarrow B_{b}$, $\mu_{0}\leftarrow\mu_{0}$, $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$, $C\leftarrow 1.01$, $K\leftarrow K$. Samples across brackets are not shared;
@@ 7      Execute $\text{alg}_{b}$ for one time round;
Algorithm 1 merges and adapts the bracketing idea of Katz-Samuels and Jamieson (2020) with the Sequential Exploration–Exploitation (SEE) framework of Li and Cheung (2025). In Katz-Samuels and Jamieson (2020), each bracket is generated by an independent random shuffle. In contrast, PSEEB defines all brackets from a single random permutation of $[K]$, yielding a nested family $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ satisfying $B_{1}\subset B_{2}\subset\cdots\subset B_{\lceil\log_{2}K\rceil+1}=[K]$. For a bracket index $1\leq b\leq\lceil\log_{2}K\rceil$, we have $|B_{b}|=2^{b}$, and $|B_{\lceil\log_{2}K\rceil+1}|=\min\{2^{\lceil\log_{2}K\rceil},K\}$. This monotone structure is crucial for our analysis, in particular for Lemma 7, which guarantees that with nontrivial probability the first several brackets contain at least one qualified arm when multiple qualified arms exist.
Lemma 7.
Given a qualified arm index $j\in[m]$, denote bracket index $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$. We have $\Pr(b_{j}\geq\tilde{b})\leq\exp\left(-\frac{j|B_{\tilde{b}-1}|}{K}\right),\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1]$. Here we denote $B_{0}=\emptyset$ and take $|B_{0}|=0$.
@@
Then, PSEEB instantiates $\lceil\log_{2}K\rceil+1$ independent algorithm copies via Algorithm 2 (given in Appendix C). This procedure is adapted from Algorithms 2–5 of Li and Cheung (2025), named as Sequential Exploration Exploitation(SEE). We adapt SEE with modified budget allocation and stopping rules to accommodate bracketing. Algorithm copy $b\in[\lceil\log_{2}K\rceil+1]$ takes bracket $B_{b}$ as an input parameter and only pull arms in $B_{b}$. Detailed discussion of Algorithm 2 is in Appendix C.
Algorithm 1 manages all the algorithm copies in a round-robin manner. In each global round (Line 1), Algorithm 1 execute one of the algorithm copies with one arm pull, and the overall algorithm terminates as soon as some copy outputs $\hat{a}\in[K]\cup\{\textsf{None}\}$. By construction, only the copy $\textnormal{alg}_{\lceil\log_{2}K\rceil+1}$ (which operates on $B_{\lceil\log_{2}K\rceil+1}=[K]$) is permitted to output None. For any other bracket $B_{b}$ with $|B_{b}|<K$, the corresponding copy never outputs None; even if it becomes convinced that $\max_{a\in B_{b}}\mu_{a}<\mu_{0}$, it has to continue sampling. This differs from SEE Li and Cheung (2025) and is necessary for the analysis of $\mathbb{E}\tau$ under bracketing.
Let $\tau$ denote the stopping time of PSEEB, and let $\tau(B_{b})$ denote the number of rounds executed by copy $\textnormal{alg}_{b}$ up to its own termination (possibly $\infty$). Since the Algorithm 1 terminates upon the first terminating copy and adopts the output, we have $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$, and hence $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{b})$.
For a positive instance $\nu$, the intuition of sample complexity $O\Big(\log K\min_{1\leq j\leq m}\big\{\log(1/\delta)/\Delta_{j,0}^{2}+H(j)\log^{4}K\log 1/\Delta_{j,0}^{2}\big\}\Big)$ in Theorem 3 is as follows. Consider a “lucky” agent with prior knowledge $j^{*}=\arg\min_{1\leq j\leq m}\{(\log(1/\delta))/\Delta_{j,0}^{2}+H(j)\}$. This agent samples $K/j^{*}$ arms in $[K]$ uniformly at random, successfully capturing an arm $j^{*}$ with reward $\mu_{j^{*}}$. However, the arm’s unknown position requires the agent $H(j^{*})$ pulls to identify $j^{*}$ at a low confidence level. Subsequently, it must pull arm $j^{*}$ another $O(\log(1/\delta)/\Delta_{j^{*},0}^{2})$ times to confirm $j^{*}$ is qualified with confidence $1-\delta$. Algorithm 1 maintains multiple parallel copies, wishing $j^{*}$ is captured in an early bracket. Since the agent cannot pre-identify the correct bracket and low-confidence identification may be erroneous, the expected sample complexity includes an additional factor of $\log^{4}K\log\frac{1}{\Delta_{j,0}^{2}}$ for the rough identification phase.
Theorems 2 and 3 are derived from several technical lemmas, leveraging Lemma 7. A proof sketch is provided in Appendix C.3, with the full proof in Appendix C.4.
4 Conclusion
This paper works on the open problem 1-identification, providing a novel lower bound for positive instances derived by a specialized optimization formulation. We further design an algorithm that is nearly optimal across all instances. Further refining logarithmic factors remains a future research.